Circular No. 8035

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jy IN NGC 477 R. Arbour, South Wonston, Hampshire, England, reports his discovery

of an apparent supernova (mag 16.3) on an unfiltered CCD image taken on Dec. 17.952 UT using a 0.3-m f/6.3 reflector. SN 2002jy is located at α = 1^{h}21^{m}16^{s}.27, δ = +40^{o}29′55′′.3 (equinox 2000.0), which is 46′′.4 west and 37′′.6 north of the center of the galaxy NGC 477. Nothing is present at this location on a 1995 blue Palomar Sky Survey plate (Digital Sky Survey). Further images taken by Arbour on Dec. 18.713 and 18.719 show the new object at about the same brightness, and an unfiltered CCD image taken by T. Boles, Coddenham, Suffolk, England, on Dec. 18.731 yields mag 16.5 and position end figures 16^{s}.23, 55′′.6. T. Vanmunster, Landen, Belgium, reports his independent discovery of SN 2002jy (providing position end figures 16^{s}.26, 55′′.5) on unfiltered CCD images taken on Dec. 18.96 with a 0.35-m f/6.3 telescope during an automated supernova-search program; a confirming CCD image was obtained on Dec. 18.991 by P. Van Cauteren and P. Lampens at Beersel, Belgium, yielding V = 16.0 for the new object. SN 2002jy was not visible on Vanmunster’s CCD image taken on Oct. 11.075 (limiting unfiltered mag 17.2) or on 1977 red Digitized Sky Survey images (limiting mag 19.5).

S/2002 J 1 Further to IAUCs 7555 and 7900, S. S. Sheppard reports the discovery

of another satellite of Jupiter with the 2.2-m reflector at Mauna Kea. The astrometry is given on MPEC 2002-Y22, along with orbital elements by B. G. Marsden showing this to be another retrograde satellite (i = 163^{o}, P = 748.5 days, H = 15.7).

V4743 SAGITTARII R. J. Ivison, Royal Observatory, Edinburgh; and T. B. Lowe, James

Clerk Maxwell Telescope, report flux densities for V4743 Sgr at various wavelengths with SCUBA, RxA3, and RxB3 detectors (0.45–1.4 mm) and with the Very Large Array (3.6 cm; uncertainties ≤ 25 percent in flux): Oct. 11.9 UT, VLA, < 0.12 mJy; 12.2, 0.45 mm, 369 ± 63; 12.2, 0.85 mm, 100 ± 10; 27.0, VLA, 0.20; 27.3, 0.85 mm, 177 ± 22; Nov. 6.2, 1.3 mm, 115 ± 23; 15.0, VLA, 1.21; 15.1, 0.9 mm, 276 ± 79; 15.1, 1.4 mm, 199 ± 57. The VLA source is located at α = 19^{h}01^{m}09^{s}.32, δ = −22^{o}00′06′′.1 (equinox 2000.0; uncertainties ≤ 0′′.5).

2002 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green